OpenStudy (anonymous):

solve using elimination method 14x +18y =-54 -4x -14y =42

OpenStudy (anonymous):

Multiplying the first equation by 2, we get 28x + 36y = - 108 Multiplying the second equation by 7, we get 28x - 98y = 294. Now solving these two equations : Adding them both we get : 28x - 28x + 36y - 98y = -108 + 294 -62y = 186 Therefore y = -186/ 62 Which is -3. Hence the answer is -3.

OpenStudy (anonymous):

thanks jas9420

OpenStudy (anonymous):

You're welcome. =]